A student walks across campus from the residence hall to the science lab in 13 minutes. The science lab is 0.8 km from the residence hall. What is the student's average speed in km/h?
1
Expert's answer
2020-09-06T17:16:04-0400
"v=\frac{s}{t}=\frac{0.8(60)}{13}=3.7\frac{km}{h}"